OpenStudy (nuccioreggie):

Simplify negative 2 over 3 ÷ negative 7 over 4. negative 7 over 6 negative 8 over 21 8 over 21 7 over 6

OpenStudy (nuccioreggie):

@paki

OpenStudy (unanimoose):

Do you mean: \[\frac{ 2 }{ 3 } \div \frac{ 7 }{ 4 }\]

OpenStudy (nuccioreggie):

@Joel_the_boss

OpenStudy (unanimoose):

If so, we need to make a common denominator.

OpenStudy (unanimoose):

We need to find the multiple of 3 and 4 that are the same so: 3, 6, 9, 12, 15 4, 8, 12, 16 There is a common multiply of 12 so we must multiply to get to that denominator and then divide.

OpenStudy (unanimoose):

\[\frac{ 2 \times 4 }{ 3 \times 4} \div \frac{ 7 \times 3}{ 4 \times 3 }\]

OpenStudy (unanimoose):

Please don't mention someone else, I'm trying to help you through. You're not just going to get a straight answer.

OpenStudy (nuccioreggie):

-2 over 3 divided by -7 over 4

OpenStudy (nuccioreggie):

i said its B

OpenStudy (unanimoose):

Then: \[\frac{ -2 \times 4 }{ 3 \times 4 } \div \frac{ -7 \times 3 }{ 4 \times 3}\]

OpenStudy (unanimoose):

I'll leave the rest to you, you can use a calculator.

OpenStudy (nuccioreggie):

WAT THE HELL ARE U TALKING ABOUT I DIDNT MENTION NOBODY ELSE WAT CHU TALKIN BOUT